Find the surface area of a cuboid (3cm by 3cm by 13cm)

Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)

(3cm *3*3*13=270
your surface area is 270

Answer by Alan3354(69443)

Find the surface area of a cuboid (3cm by 3cm by 13cm)
----------
SA = 2*(3*3 + 3*13 + 3*13)
= 2*(9 + 39 + 39)
= 174 sq cm